OpenStudy (anonymous):

dose anyone have a kid on here? cuz i want you all to work with me its hard trying to be a single dad and be in high school!!

OpenStudy (anonymous):

whats the problem?

OpenStudy (anonymous):

(1/4)A=23

OpenStudy (anonymous):

that means one-fourth of a number , lets call it A, is 23. so if one fourth of A is 23, what would A be?

OpenStudy (anonymous):

1/3??

OpenStudy (anonymous):

because you would devied 23 by one forth right?

OpenStudy (jamesj):

No, 1/3 is not right. Suppose (1/4)A = 1. What would A be in this case?

OpenStudy (jamesj):

I.e., suppose \[\frac{A}{4} = 1\] What is A?